OpenStudy (anonymous):

simplify 4x/9 + x-10/9 ?

OpenStudy (anonymous):

4x/9 + x - 10/9 4x/9 + 9x/9 - 10/9 13x/9 - 10/9 or if you want to put everything over the common denominator, 9, write it like below: \[\frac{ 13x-10}{9}\]

OpenStudy (anonymous):

Im sorry but that equation came out wrong. This is the legit one.

OpenStudy (anonymous):

\[\frac{ 4x }{9} - \frac{ x-10 }{ 9 } = \frac{ 4x - (x-10) }{ 9 } = \frac{ 4x-x+10 }{ 9 }\] \[\frac{ 3x-10 }{ 9 } = \frac{ 3x }{ 9 } - \frac{ 10 }{ 9 } = \frac{ x }{ 3 } - \frac{ 10 }{ 9 }\]

OpenStudy (anonymous):

Bro! Thank you sooo much. Could you please help me with this on too

OpenStudy (anonymous):

Combine the terms in the numerator... \[\frac{ t ^{2}+6t - 7 }{ t-1 }\] Often the next step is to factor the numerator polynomial... you hope to be able to factor out the same term that is in the denominator, in this case (t-1)

OpenStudy (anonymous):

\[\frac{ t ^{2} +6t - 7}{ t-1 } = \frac{ (t-1)(t+7) }{ t-1 }\] then cancel the (t-1) terms in numerator and denominator to be left with t+7